Vanessa built a rectangular pen for her dogs. she used an outside wall of the garage for one side of the pen. she used 20 m of fencing in order to build the other 3 sides to complete the rectangle. find the dimensions of the pen of the pen if the area of the resulting pen was 48 m^2
Answer by mananth(16949)

Vanessa built a rectangular pen for her dogs. she used an outside wall of the garage for one side of the pen. she used 20 m of fencing in order to build the other 3 sides to complete the rectangle. find the dimensions of the pen of the pen if the area of the resulting pen was 48 m^2
Let one side be x
the other side = 48/x
2x+48/x =20
2x^2+48=20x
2x^2-20x+48=0
x^2-10x+24=0
x^2-6x-4x+24=0
x(x-6)-4(x-6)=0
(x-6)(x-4)=0
x=6 or 4 x=6 since we do not know whcich side of the garage was used
the dimensions of the pen could be 6m & 8m or 4 m & 12 m